OpenStudy (anonymous):

slope-intercept form .... slope=.25, x-intercept=24

OpenStudy (aroub):

same thing!

OpenStudy (anonymous):

y=.25+24 ??

OpenStudy (anonymous):

y=.25x+24

OpenStudy (anonymous):

if that is the x-intercept then that is not correct. if it is the y-intercept, it is correct!

OpenStudy (anonymous):

i think im wrong what do i do all i given is the slope.25 and the xintecept =24

OpenStudy (anonymous):

x-intercept is (24,0) y=mx+b 0=.25(24)+b 0=6+b b=-6 y=0.25x-6